OpenStudy (anonymous):

Use the compound interest formula, A = P times the quantity of 1 + r over n where the quantity is to the n t power. , to find the following amount of money after the given amount of time: Alexia has $4864.77 in her money market account currently. If the annual interest rate is 3.9%, and the interest is compounded twice per year, how much money will be in her account after 2 years?

OpenStudy (texaschic101):

A = P ( 1 + r/n) ^nt A = 4864.77 (1 + .039/2)^(2)(2) A = 4864.77 ( 1 + .045) ^ 4 A = 4864.77 (1.045) ^ 4 A = 4864.77 (1.1925) A = 5801.24